# Good News for Polynomial Root-finding

### v35

@@ 1.6.3 Cauchy-based root-counting and e/i tests
In Secs. 1.6.4 and 1.6.5 we outline our two approaches to devising e/i tests for a black box polynomial based on (1.9) for $\mathcal{D}=D(0,1)$ We cover these two approaches in some detail in Part II, made up of Secs. 29, and Part III, made up of Secs. 10-15, respectively.
The subdivision root-finders that incorporate these e/i tests are quite fast but only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$. In Part IV, made up of Secs. 16-20 and outlined in Sec. 1.6.7, we fix both deficiencies, thus proving our Main Theorems.
1.6.4 Our soft e/i tests in Part II: outline
@@
(i) choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$,
(ii) fix a sufficiently large integer $q$ of order $m\log(d)$, and
(iii) compute approximation $s_{0,q}$ to the integral (1.9) for the disc $D(0,\rho)$ and the polynomial $t(x)=p(\rho x)$ replacing $p(x)$.
@@
(iv) Obtain #$\#(D(0,\rho))$ as the closest integer approximation of $s_{0,q}$, prove that the recipe is correct whp, and extend the root-counter to e/i test.
(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$.
The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points; in Part II we also devise a deterministic soft e/i test at the price of increasing the above bound on the number of points to order of $m^{2}\log(d)$.
1.6.5 Our generic soft e/i test in Part III
(i) In Part III we decrease our estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer vanishes or not, which is sufficient for e/i tests. In these e/i tests we elaborate upon the following generic one, where we use $h$-lifting of roots generalizing DLG root-squaring (1.5):
@@ 1.6.5 Our generic soft e/i test in Part III
for $h=1,2,\dots$ and $t_{2^{k}}(x)=p_{k}(x)$ for all $k$.
@@
INPUT: a black box polynomial $p(x)$ of a degree $d$, given by an oracle for its evaluation, a real $\sigma>1$ and a positive integer $h$ (both specified in Sec. 11).
INITIALIZATION: Fix a positive $q\in\mathbb{N}$.
@@
(2) By extending (1.6) to $p_{h}(x)$ of (1.5) (we specify this extension in (10.5)) closely approximate
$v_{i}:=|{\rm NIR}_{p_{h}}(x_{i})|=\Big{|}\frac{p^{\prime}_{h}(x_{i})}{p_{h}(x_{i})}\Big{|},~{}i=1,\dots,q.$ (1.12)
(3) Certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion, if $1+d/v_{j}\leq\sigma^{h}$ for some $i\leq q$. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
(ii) To prove correctness of soft inclusion claim of Alg. 1.1, first differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$ and obtain the following well-known equation:
${\rm NIR}(x):=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}}.$ (1.13)
@@
#$\#(D(c,\rho))>0$ if $\rho\leq d~{}|p(c)/p^{\prime}(c)|$.
@@
We use Cor. 1.3 later but next apply Cor. 1.4 to prove the following lemma, which implies correctness of the soft inclusion claim of Alg. 1.1.
@@
Apply Cor. 1.4 to the polynomial $p_{h}(x)$ and obtain that the equation $\rho=d\Big{|}\frac{p_{h}(x_{)}}{p^{\prime}_{h}(x)}\Big{|}$ implies that the disc $D(x,\rho)$ contains a zero of $p_{h}(x)$. Hence the disc $D(0,|x|+\rho)$ contains a zero of $p_{h}(x)$ by virtue of triangle inequality. Therefore, the disc $D(0,\sigma)$ contains a zero of $p(x)$ for $\sigma=(|x|+\rho)^{1/h}$. ∎
Apply this lemma for $x=x_{i}$, $|x_{i}|=1$, $\rho=d/v_{i}$, and $v_{i}=|{\rm NIR}_{p_{h}}(x_{i})|$ (cf. Alg. 1.1). Deduce that #$(D(0,\sigma))>0$ for $\sigma^{h}=1+d/v_{i}$, that is, Alg. 1.1 is correct when it claims $\sigma$-soft inclusion.
Observation 1.4.
$\sigma$-soft inclusion test by means of Alg. 1.1 is correct where
@@
Apply Lemma 1.1 and obtain that #$(D(0,\sigma))>0$ for $\sigma=(1+\rho)^{1/h}=2^{\nu}~{}{\rm for}~{}\nu=\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}$.
@@
(b) We prove correctness of the exclusion claim of Alg. 1.1 separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (1.5) or root-lifting (1.11). In Sec. 11.3 we choose any integer $q$ exceeding $d$ and then prove deterministic correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity. This bound is compatible with the bound of Example 1.4 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$. We also extend this deterministic e/i test by allowing any $q>m$ (rather than $q>d$) provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
@@
In Sec. 11.4 we sample $q=O(\log(d))$ iid random values of $x_{i}$ under the uniform probability distribution on the circle $C(0,1)$ and then prove that the exclusion claim is correct whp under the Random Root Model if $|{{\rm NIR}_{p_{h}}(x_{i})}|\leq 1/d^{a}$ for a sufficiently large constant $a$.
@@
(iv) Counting query locations. Approximation of $p_{h}(x)$ is reduced to approximation of $p(x)$ at all the $h$th roots of $x$ for $h$ being a power of 2 based on (1.5) or for any integer $h>1$ if we apply (1.11), and so Alg. 1.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$, when it claims inclusion, or at $2qh$ points, when it claims exclusion. In all these estimates we choose factor $h$ in $O(\log(d))$, and so we can ignore this factor in our $\tilde{O}$ estimates and say that our soft e/i test by means of Alg. 1.1 evaluates polynomial $p(x)$ at $\tilde{O}(m)$ points. Combine these estimates with Cor. 1.2 to bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
@@
Under the Random Root Model of Sec. 1.5 the zeros of $p$ are isolated whp from the unit circle $C(0,1)$, and we decrease the bound $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$ whp by exploit their even stronger isolation ensured by means of $O(\log(\log(d)))$ root-squaring steps (1.5) or alternatively with $O(\log(d))$ steps of root-lifting of (1.11).
1.6.6 Computational precision and bit-operation complexity
@@
Louis and Vempala in [73] bound by $O(b\log(d))$ the computational precision of their novel high order Newton’s iterations. We also prove this bound for our root-finders by using a small part of their analysis together with Las Vegas randomization and some rather customary techniques of numerical analysis.
We extend this bound combined with the bound on the number of queries in our root-finders to the bit operation complexity estimates: we use the known fast algorithms for multipoint polynomial evaluation. We satisfy the assumptions of the latter algorithms by using map (1.4).
@@
We devise two groups of supporting algorithms in Parts II and III of the paper, and in both cases arrive at the estimates for the numbers of evaluation points and bit-operations involved in our algorithms that exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
1.6.7 Acceleration of root-finding in Part IV: outline
@@
In Part IV we accelerate our algorithms to yield the Main Theorems. Here is an outline of that acceleration.
@@
Represent all partitions of components with a component tree as follows. At most $m$ components, said to be the leaves, are covered by isolated $\epsilon$-discs for a fixed error tolerance $\epsilon\geq 0$. The other (at most $m-1$) vertices of the tree are components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree. The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between their partitions.
Recall that every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step. Hence in $O(\log_{2}(mb))$ steps, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
@@
(ii) Acceleration of subdivision process with compression of components. The bound $O(m)$ does not generally hold for the overall number of suspect squares at all subdivision steps. This is because subdivision steps become inefficient where a component stays compact and not broken in many consecutive subdivision steps (see Fig. 3), but we replace these steps with compression of such a component $C$ into a near minimal disc $D$ that shares with $C$ all the zeros of $p$ lying in it.
@@
If the disc $D$ is an $\epsilon$-disc we just output it. Otherwise we apply subdivision iterations to its minimal superscribing square $S$. Since we compress the component into a near minimal disc $D$, we obtain readily that these iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc. From this we deduce that our subdivision process with compression reduces the solution of Problem 1 to at most $22m-$ 2 compression steps and $O(m)$ soft e/i steps overall.
(iii) Compression of components. We reduce a compression of a compact component to the following main stages (see more details in Sec. 20.5):
(i) compute a point $c$ lying in or near the minimal disc that covers all zeros of $p$ lying in the component $C$ and let $c$ denote the center of the compressed disc that we seek;
(ii) compute a close upper bound $\rho$ on the distance from $c$ to the farthest zero of $p$ among its at most $m$ zeros lying in the compact component $C$.
We prove that Schröder’s single celebrated iteration (20.1) following $O(\log(d))$ auxiliary subdivision steps performs step (i) above, that is, computes the desired center $c$.
Beginning stage (ii) above we have a range $[\rho_{-},\rho_{+}]$ where the bound $\rho$ lies, for $10<\rho_{-}<\rho_{+}<$ 1. Hence we can approximate it in $\lceil\log_{2}(1/\epsilon)\rceil$ bisection steps, but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the exponents, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]$.
We extend our soft e/i tests to define the pivot step of such a modified bisection. Summarizing we reduce the solution of Problem 1 to $\tilde{O}(m)$ soft e/i tests and $\tilde{O}(m)$ Schröder’s iterations.
In this way we accelerate the root-finders of Parts II and III by a factor of $b$. Together with our recipes for detection and correction of output errors of root-finders in the next subsection, this completes the proof of our Main Theorems.
@@
For bounding precision of computing, Louis and Vempala in [73] presented a major technical innovation, which enabled us to deduce a priori estimates for the bit operation cost of our subdivision root-finders versus routine a posteriori estimates in the previous papers on subdivision root-finding (see, e.g., [118, 23]). In particular, the bit operation cost estimates of the Main Theorem of [23] contain the term
@@
1.7 Error detection and correction
Proposition 1.2.
Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we can avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (1.13):
In many cases one can accelerate simultaneous numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ at many points $x$ by applying Fast Multipole Method (FMM) of [48, 27, 30].
1.8 Our novelties and related works
1.8.1 Our novelties (a short list)

### v36

@@ 3.3.1 Introductory comments
3.3.2 Cauchy-based soft e/i tests: outline
@@
(i) Choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$.
(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
@@
(iv) Conclude that exclusion holds whp if $|s_{0,q}|<1/2$.
(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$, from $\frac{1}{\gamma}$ to $\frac{1}{\gamma^{v}}$ for $\gamma>1$ of our choice, say, for $\gamma=2$.
The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points.
3.4 Soft e/i test with root-lifting: generic algorithm
In Part III we decrease the complexity estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer is 0 or not, which is sufficient for e/i tests. In these e/i tests we elaborate upon the generic one (see Alg. 3.1 below), where we use $h$-lifting of roots:
@@ 3.4 Soft e/i test with root-lifting: generic algorithm
for $h=1,2,\dots$ and $t_{2^{k}}(x)=p_{k}(x)$ for all $k$. $h$-lifting maps roots to their $h$th powers; this keeps the unit disc $D(0,1)$ and its boundary circle $C(0,1)$ intact but moves the external roots away from the circle, thus strengthening its isolation from those roots.
Remark 3.1.
The $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [51]):
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i}}(x^{2}):=(-1)^{d}p_{2^{i-1}}(x)p_{2^{i-1}}(-x)=\prod_{j=1}^{d}(x-z_{j}^{2^{i}}),~{}i=1,2,\dots.$ (3.7)
Polynomial root-finding in the 19-th and well into the 20-th century relied on iterations (3.7) combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots. The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.3) from any complex point (see [86, 113, 115] for details). The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials. Extensive tests with standard test polynomials have consistently demonstrated efficiency of the particular implementation of this approach in [115], but even the straightforward implementation of iterations (3.7) promises to be highly efficient as well.
We will elaborate upon the following generic algorithm in Sec. 14.
@@
INPUT: $\sigma>1$, an integer $d>0$, and a $d$-th degree black box polynomial $p(x)$ given by an oracle for its evaluation.
INITIALIZATION: Fix a positive $q\in\mathbb{N}$ and a positive integer $h$.
@@
(2) By extending (3.2) to $p_{h}(x)$ of (3.7) (we specify this extension in (13.3)) approximate the values
$v_{g}:=|{\rm NIR}_{p_{h}}(x_{g})|=\Big{|}\frac{p^{\prime}_{h}(x_{g})}{p_{h}(x_{g})}\Big{|},~{}g=1,\dots,q.$ (3.8)
(3) If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
3.5 Soft e/i test with root-lifting: correctness of inclusion claim
@@
It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D(c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
@@
We use Cor. 3.1 later but next apply Cor. 3.2 to prove correctness of the inclusion claim of Alg. 3.1. The bound of Cor. 3.1 on $r_{d}(c)$ is too large for that, but a similar bound on $r_{d}(c,p_{h})=(r_{d}(c))^{h}$, implies that $r_{d}(c)$ is close to 1 already for $h$ of order $\log(d)$.
@@
Apply Cor. 3.2 to the polynomial $p_{h}(x)$ and obtain that the equation $\rho=d\Big{|}\frac{p_{h}(x_{)}}{p^{\prime}_{h}(x)}\Big{|}$ implies that the disc $D(x,\rho)$ contains a zero of $p_{h}(x)$. Hence the disc $D(0,|x|+\rho)$ contains a zero of $p_{h}(x)$ by virtue of triangle inequality. Therefore, the disc $D(0,\sigma)$ contains a zero of $p(x)$ for $\sigma=(|x|+\rho)^{1/h}$. ∎
Apply this lemma for $x=x_{i}$, $|x_{i}|=1$, $\rho=d/v_{i}$, and $v_{i}=|{\rm NIR}_{p_{h}}(x_{i})|$ (cf. Alg. 3.1). Deduce that #$(D(0,\sigma))>0$ for $\sigma^{h}=1+d/v_{i}$, that is, Alg. 3.1 is correct when it claims $\sigma$-soft inclusion.
Observation 3.1.
$\sigma$-soft inclusion test by means of Alg. 3.1 is correct where
@@
Apply Lemma 3.1 and obtain that #$(D(0,\sigma))>0$ for $\sigma=(1+\rho)^{1/h}=2^{\nu}~{}{\rm and}~{}\nu=\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}$.
@@
3.6 Soft e/i test with root-lifting: further analysis (outline)
@@
(i) Correctness of the exclusion claim of Alg. 3.1. We prove it separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (3.7) or root-lifting (3.6).
(a) In Sec. 14.3 we choose any integer $q$ exceeding $d$ and then prove deterministic correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity. This bound is compatible with the bound of Example 3.1 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$. We also extend this deterministic e/i test by allowing any $q>m$ (rather than $q>d$) provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
@@
(b) In Sec. 14.4 we sample $q=O(\log(d))$ iid random values of $x_{i}$ under the uniform probability distribution on the circle $C(0,1)$ and then prove that the exclusion claim is correct whp under the Random Root Model if $|{{\rm NIR}_{p_{h}}(x_{i})}|\leq 1/d^{a}$ for a sufficiently large constant $a$.
@@
(ii) Counting query locations. Eqn. (3.7) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all the $h$th roots of $c$. We can fix $h>1$ being a power of 2 and apply (3.7) or fix any integer $h>1$ and apply (3.6). In both cases Alg. 3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$, when it claims inclusion, or at $2qh$ points, when it claims exclusion. We always choose $h=O(\log(d))$ and then can ignore a factor of $h$ in our $\tilde{O}$ estimates.
In our deterministic soft e/i test based on Alg. 3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points, combine these estimates with Cor. 2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
@@
Under the Random Root Model of Sec. 1.5 the roots are reasonably well isolated whp from the unit circle $C(0,1)$. We strengthen this isolation by applying $O(\log(\log(d)))$ root-squaring steps (3.7) or $O(\log(d))$ steps of root-lifting of (3.6) and then whp decrease the overall bit operation count from $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$.
@@
(iii) Estimation of computational precision and bit-operation complexity. Louis and Vempala in [76] bound by $O(b\log(d))$ the computational precision of their novel high order Newton’s iterations. We combine a small part of their advanced analysis with Las Vegas randomization and some customary techniques of error analysis to extend this estimate to our root-finders.
For the solution of Problems 0, 1, and 3, where the coefficients are given, we apply fast algorithm of [81] for multipoint polynomial evaluation to decrease the estimated number of queries and the bit operation cost bound.
@@
We devise two groups of efficient e/i tests in Parts II and III of this paper, and in both cases arrive at the estimates for the numbers of (i) evaluation points, that is, of the oracle queries, and (ii) bit-operations involved in our algorithms. Both of the estimates exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
4 Acceleration of subdivision root-finding by means of compression: outline
@@
In Part IV we accelerate the algorithms of both Parts II and III by these factors (of $b$ for roots and $\tilde{b}$ for eigenvalues) to prove the Main Theorems in two ways. Next we outline that acceleration.
@@
Represent all partitions of components with a component tree, $\mathbb{T}$, as follows. At most $m$ components, said to be the leaves, are covered by isolated $\epsilon$-discs for a fixed error tolerance $\epsilon\geq 0$. The other (at most $m-1$) vertices of the tree are components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree. The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between their partitions.
Recall that every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step. Hence in $O(\log(b))$ subdivision iterations, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
@@
(ii) Acceleration of subdivision process with compression of components. Subdivision process becomes quite inefficient where a component stays compact and not broken into more than one component in many consecutive subdivision steps (see Fig. 3). To accelerate it, we replace these steps with compression of such a component into an $\epsilon$-disc or a near-minimal disc $D$ that still keeps all roots lying in the component. In Sec. 19 we comment on two customary but inferior alternatives to our compression, relying on deflation or solely on functional iterations.
@@
If the compression outputs $\epsilon$-disc, we just output it and do not subdivide it anymore. Otherwise we apply subdivision iterations to the minimal superscribing square $S$ of the disc $D$. Since the disc is near-minimal, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc. From this we deduce that our subdivision process with compression reduces the solution of Problem 1 to at most $22m-$ 2 compression steps and $O(m)$ soft e/i steps overall.
(iii) Compression of components. We reduce compression of a compact component $C$ to the following two main stages (see more details in Sec. 23.5):
(a) compute a point $c^{\prime}$ lying in or near the minimal disc that covers all roots lying in the component $C$ and let $c^{\prime}$ denote the center of the compressed disc that we seek;
(b) compute a close upper bound $\rho^{\prime}$ on the maximal distance from $c^{\prime}$ to the roots in the component $C$ and output a compressed disc $D(c^{\prime},\rho^{\prime})$.
We prove that Schröder’s single iteration (23.1) following $O(\log(d))$ auxiliary subdivision steps performs step (a) above, that is, computes the desired center $c^{\prime}$.
At the beginning of step (b) above we have a range $[\rho_{-},\rho_{+}]$ for the bound $\rho^{\prime}$, $0<\epsilon\leq\rho_{-}<\rho_{+}\leq 1$. Hence we can approximate it in $\lceil b\rceil$ bisection steps, for $b=\log_{2}(1/\epsilon)$ but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the logarithms, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
We extend our fast soft e/i tests of Part II to fast soft $m$-tests (see part (A) of Sec. 2.1) and use them to define the pivot step of such a modified bisection.
Summarizing we reduce the solution of Problems 0 and 1 to $\tilde{O}(m)$ soft e/i tests and $m$ tests and $\tilde{O}(m)$ Schröder’s iterations. In this way we accelerate the root-finders of both Parts II and III by a factor of $b/\log(b)$ to complete the proof of our Main Theorems.
@@
By using some of the advanced techniques of Louis and Vempala in [76], we readily bound the precision of computing in our root-finders. This enables us to deduce a priori estimates for the bit operation cost of our subdivision root-finders versus inferior known a posteriori estimates (see, e.g., [122, 24]). In particular, the bit operation cost estimates of the Main Theorem of [24] contain the term
@@
5 Our novelties, related works, implementation, and further progress
5.1 Our novelties (a short list)